OpenStudy (anonymous):

Find y'' in terms of x and y by implicit differentiation. x^4+y^4=3^4 working on this for a while please help

OpenStudy (abb0t):

take the derivative as you would

OpenStudy (anonymous):

\[4x^3+4y^3(\frac{ dy }{ dx })=0\]\[\frac{ dy }{ dx }=- \frac{ x^3 }{ y^3 }\]\[12x^2+12y^2(\frac{ dy }{ dx })+\frac{ d^2y }{ dx^2 }(4y^2)=0\]\[\frac{ d^2y }{ dx^2 }=\frac{ x^3 }{ 48y^3 }-\frac{ 3x^2 }{ y^2}=\frac{ x^3-144x^2y }{ 48y^3 }\]

OpenStudy (anonymous):

Sorry, I made a mistake in the above math in the second step. This is the second derivative now that I have fixed my mistake: \[\frac{ d^2y }{ dx^2 }=\frac{ x^6-144x^2y^4 }{ 48y^7 }\] @abb0t

OpenStudy (anonymous):

@Solongos